### v1

# Impact of the valence band on Rydberg excitons in cuprous oxide quantum wells

Abstract
The complex valence band structure of bulk cuprous oxide necessitates going beyond the parabolic approximation to precisely estimate exciton binding energies. The same is true for excitons in cuprous oxide quantum wells, for which many effects have been obtained so far only qualitatively within a hydrogenlike two-band model. Here, we derive the complete Hamiltonian for excitons in cuprous oxide quantum wells based on the Luttinger-Kohn model, taking into account the full complex valence band structure. Symmetry properties of the system are discussed. Numerical results based on the diagonalization of the Hamiltonian using B-spline functions reveal the energy shifts and the lifting of degeneracies due to the nondiagonal coupling terms of the complex valence band. The relative oscillator strengths of the excitonic transitions induced by circularly polarized light are also calculated.
I Introduction
Electron-hole bound states in bulk cuprous oxide are characterized by large binding energies, allowing for high-resolution measurements of the excitonic transitions of yellow and green Rydberg series [1, 2, 3]. This paves the way for excitonic devices, sensors [4, 5] and potentially room-temperature applications [6, 7, 8]. Due to strong spin-orbit coupling, an accurate theoretical description of exciton spectra in bulk Cu2O requires taking into account the complex valence band structure that goes beyond the simple parabolic band approximation. Effects of the band structure on exciton energies have been investigated recently both theoretically and experimentally [9, 10, 11, 12, 13, 14, 15]. These works have significantly advanced our understanding of the exciton properties in bulk Cu2O, in comparison to the pioneering studies of exciton energies and absorption [16, 17, 18].
Similar to other types of semiconductors such as GaAs [19] or CdSe [20, 21], a greater tunability of electronic and optical properties of quasiparticles in Cu2O can be achieved by fabrication of heterostructures [22, 23]. The first films of cuprous oxide sandwiched between sapphire substrates and thus realizing the weak quantum confinement of lower exciton states have already been obtained [24, 25]. When the size of a particular exciton state is comparable to the thickness of the film, the confinement effect dominates. Therefore, for very thin Cu2O films or for Rydberg exciton states, the regime of strong confinement is reached, and the details of the complex band structure play the role of a small perturbation. As a result, the parabolic approximation of the bands in the vicinity of the $\Gamma$-point is a reasonable choice. However, for weak confinement the effects due to features of the band structure are comparable to the quantum-confinement effects. Thus, the full Cu2O band structure has to be explicitly taken into account.
The quantum confinement appearing due to the band offsets at the heterojunction can, in its simplest form, be modeled by the quantum well (QW) potential. Excitons in QWs exhibit a large variety of interesting phenomena and effects [26, 27]. The theory of excitons and their radiative properties in GaAs-based QWs is well developed [19], though for cuprous oxide finite-sized crystals the energy spectrum as well as the relative oscillator strengths have been obtained only very recently [28, 29]. These works significantly differ from a simplified consideration in Refs. [30, 31] by studying the spectrum as a function of the QW width. A dependence of the exciton energy levels on the width of a QW allows one to study a crossover from weak confinement in the bulk to nearly two-dimensional excitons in case of strong confinement. Besides, above the scattering threshold electron-hole resonances occur. Their spectral lines are usually additionally broadened due to coupling to the continuum. For GaAs-based QWs, the energies and linewidths of the electron-hole resonances have been estimated using Feshbach theory [32] in Refs. [33, 34]. The resonant broadenings can also be computed by both the stabilization and the complex-coordinate-rotation method, as it was done for GaAs-based [35] and Cu2O-based QWs [36]. When varying the confinement strength, so-called bound states in the continuum (BICs), i.e., long-lived Rydberg states of the confined electron-hole pairs, appear in the continuum background [37]. The BICs are uncoupled from the continuum of lower quantum-confinement subbands and thus have zero linewidths [38, 39]. All these effects can be successfully studied numerically by expanding the electron-hole wave function over a basis of B-splines [40, 41]. The wave functions in coordinate space, computed using a B-spline representation, clearly illustrate, e.g., the quenching of states in the strong confinement regime, the localized nature of BICs and, furthermore, allow for the estimation of oscillator strengths.
In Refs. [36, 28, 29], the energy levels of excitons and electron-hole resonances in finite-sized cuprous oxide crystals have been calculated. Moreover, these works shed light on the crossover from weak to strong confinement regime as well as the radiative properties of exciton states of different symmetries. However, they took into account only the effect of the dielectric contrast in the crystal and in the substrate [28] which, for thin films of semiconductors, is usually described by the Rytova-Keldysh potential [42, 43]. The nonparabolicity of the valence bands near the $\Gamma$-point has not been considered so far. Taking into account the nonparabolic energy dispersion of the hole leads to an increase in the number of coupled differential equations originating from the complex valence band. This significantly complicates the numerical treatment of the problem. Fortunately, the B-spline expansion of the exciton wave function still allows for an effective numerical solution of this challenging problem.
The aim of this article is to study the impact of the complex valence band on excitons in cuprous oxide QWs. Besides the quantization over the confinement direction, we consider the Luttinger Hamiltonian [44] or, equivalently, the Suzuki-Hensel Hamiltonian [45] for the hole dispersion in the QW together with a purely parabolic dispersion for the electron. Such a hole dispersion significantly complicates the Schrödinger equation for excitons in cuprous oxide QWs [28, 29]. Moreover, it breaks some exact symmetries of the pure parabolic dispersion, in particular, the rotational invariance around the axis normal to the QW plane, making the angular momentum quantum number $m$ to be only an approximate quantum number. Nevertheless, using the B-spline representation of the wave function, we can calculate the matrix elements of all involved operators and solve the eigenvalue problem to obtain the energy levels as a function of the QW width. We study effects of the complex valence band, originating from coupling of the states with different values of $m$. We also derive expressions for the relative oscillator strengths of the excitonic transitions induced by circularly polarized light. We show that the oscillator strengths are related to derivatives of certain projections of the exciton wave function at zero separation between electron and hole.
The article is organized as follows. In Sec. II, we derive the theory for excitons in cuprous oxide QWs taking into account the complex valence band structure. We obtain the Hamiltonian, expressions for relative oscillator strengths and provide a numerical algorithm for the computation of exciton spectra and wavefunctions. Results are presented and discussed in Sec. III. Concluding remarks are given in Sec. IV.
II Theory and methods
In this section, we describe excitons in cuprous oxide QWs by taking into account the complex structure of the valence band. The general setup of our system is illustrated in Fig. 1.
Figure 1: Sketch of the general setup for the creation of excitons in a cuprous oxide QW. The $[001]$ axis of the crystal is aligned perpendicular to the QW plane. Excitons are excited by circularly polarized light along this axis.
Let $L$ be the width of a thin layer of cuprous oxide sandwiched between another material with larger band gap. Due to band offsets at the heterojunction, this system can be regarded as a QW. Excitons in the QW are created by circularly polarized laser light along the normal direction to the QW plane. In general, the exciton spectra depend on the orientation of the crystal relative to the QW. In this article, we consider a geometry in which the principal axis $[001]$ of the crystal is aligned along the $z$ direction perpendicular to the QW plane. The Hamiltonian including the impact of the valence band is derived in Sec. II.1 and its symmetry properties are discussed in Sec. II.2. The computation of the oscillator strengths and the numerical algorithm for the diagonalization of the Hamiltonian using an appropriate basis set are explained in Secs. II.3 and II.4, respectively.
II.1 Hamiltonian
The Hamiltonian for electron and hole in the QW is given by
$H$ $=H_{\mathrm{e}}(\boldsymbol{p}_{\mathrm{e}})+H_{\mathrm{h}}(\boldsymbol{p}_{\mathrm{h}})-\frac{e^{2}}{4\pi\varepsilon_{0}\varepsilon}\frac{1}{|\boldsymbol{r}_{\mathrm{e}}-\boldsymbol{r}_{\mathrm{h}}|}$
$+V_{\mathrm{e}}(z_{\mathrm{e}})+V_{\mathrm{h}}(z_{\mathrm{h}})$ (1)
with
$H_{\mathrm{e}}(\boldsymbol{p}_{\mathrm{e}})$ $=E_{\mathrm{g}}+\frac{\boldsymbol{p}_{\mathrm{e}}^{2}}{2m_{\mathrm{e}}}\,$, (2)
$H_{\mathrm{h}}(\boldsymbol{p}_{\mathrm{h}})$ $=H_{\mathrm{SO}}+\frac{1}{2\hbar^{2}m_{0}}\Big[\hbar^{2}\left(\gamma_{1}+4\gamma_{2}\right)\boldsymbol{p}_{\mathrm{h}}^{2}$
$+2\left(\eta_{1}+2\eta_{2}\right)\boldsymbol{p}_{\mathrm{h}}^{2}\left(\boldsymbol{I}\cdot\boldsymbol{S}_{\mathrm{h}}\right)-6\gamma_{2}\left(p_{1\mathrm{h}}^{2}\boldsymbol{I}_{1}^{2}+\mathrm{c.p.}\right)$
$-12\eta_{2}\left(p_{1\mathrm{h}}^{2}\boldsymbol{I}_{1}\boldsymbol{S}_{\mathrm{h}1}+\mathrm{c.p.}\right)$
$-12\gamma_{3}\left[\left\{p_{1\mathrm{h}},p_{2\mathrm{h}}\right\}\left\{\boldsymbol{I}_{1},\boldsymbol{I}_{2}\right\}+\mathrm{c.p.}\right]$
$-12\eta_{3}\left[\left\{p_{1{\mathrm{h}}},p_{2{\mathrm{h}}}\right\}\left(\boldsymbol{I}_{1}\boldsymbol{S}_{\mathrm{h}2}+\boldsymbol{I}_{2}\boldsymbol{S}_{\mathrm{h}1}\right)+\mathrm{c.p.}\right]\Big]\,$, (3)
$H_{\mathrm{SO}}$ $=\frac{2}{3}\Delta\left(1+\frac{1}{\hbar^{2}}\boldsymbol{I}\cdot\boldsymbol{S}_{\mathrm{h}}\right)\,.$ (4)
Here, $E_{\mathrm{g}}$ is the gap energy and $\gamma_{i},\eta_{i}$ are the Luttinger parameters.
Figure 2: Sketch of the band structure of the $\Gamma_{6}^{+}$ conduction bands, as well as the $\Gamma_{7}^{+}$ and $\Gamma_{8}^{+}$ valence bands near the $\Gamma$ point.
The inclusion of the full set of Luttinger parameters leads to the non-trivial band structure shown in Fig. 2. Using the notation for the symmetry group $O_{\mathrm{h}}$ [46], we can classify the bands as in the bulk crystal. The lowest conduction band has the symmetry $\Gamma_{6}^{+}$, whereas the uppermost valence bands are split into the uppermost $\Gamma_{7}^{+}$ bands associated with the yellow exciton Rydberg series, and the lower $\Gamma_{8}^{+}$ bands associated with the green series. Note that for excitons in the QW we use the same values of the Luttinger parameters as in the bulk. The material parameters of cuprous oxide are given in Table 1. $\boldsymbol{I}$ and $\boldsymbol{S}_{\mathrm{h}}$ are the quasispin with $I=1$ and hole spin with $S_{\mathrm{h}}=1/2$, whose components fulfill the commutation relations of an angular momentum or spin, $\Delta$ is the spin-orbit coupling parameter, $\{a,b\}=(ab+ba)/2$ is the symmetrized product, and c.p. denotes cyclic permutation.
Table 1: Material parameters of Cu2O used in the calculations.
Energy gap $E_{\rm g}=2.17208\,$eV [1]
Spin-orbit coupling $\Delta=0.131\,$eV [10]
Effective electron mass $m_{\rm e}=0.99m_{0}$ [47]
Dielectric constant $\varepsilon=7.5$ [48]
Valence band parameters $\gamma_{1}=1.76$ [10]
$\gamma_{2}=0.7532$ [10]
$\gamma_{3}=-0.3668$ [10]
$\eta_{1}=-0.020$ [10]
$\eta_{2}=-0.0037$ [10]
$\eta_{3}=-0.0337$ [10]
The potentials $V_{\mathrm{e},\mathrm{h}}(z_{\mathrm{e},\mathrm{h}})$ confine the motion of electron and hole across the QW along the $z$ direction. For a finite-sized crystal surrounded by vacuum or air, the confinement potential of the QW can be well approximated by an infinite potential well,
$V_{\mathrm{e,h}}(z_{\mathrm{e,h}})=\left\{\begin{array}[]{lr}0&\mbox{ if }|z_{\mathrm{e,h}}|<L/2\\ \infty&\mbox{ if }|z_{\mathrm{e,h}}|\geq L/2\end{array}\right.$. (5)
The energies of electron and hole states in such a QW,
$E_{N_{\mathrm{e}},N_{\mathrm{h}}}(L)=\frac{\hbar^{2}\pi^{2}}{2L^{2}}\left(\frac{N_{\mathrm{e}}^{2}}{m_{\mathrm{e}}}+\frac{N_{\mathrm{h}}^{2}}{m_{\mathrm{h}}}\right),$ (6)
can be described by two quantum numbers $N_{\mathrm{e}},N_{\mathrm{h}}=1,2,3,\dots$. These quantum-confinement energies play the role of scattering thresholds, where bound states or resonances can accumulate [36]. Note that we use the same dielectric constant $\varepsilon=7.5$ for the crystal and the substrate and thus do not consider the dielectric contrast described by the Rytova-Keldysh potential [42, 43] for the interaction between electron and hole as discussed, e.g., in Ref. [28]. We also do not consider central-cell corrections, which can affect even-parity excitons [49].
In contrast to the bulk with its translational invariance in three dimensions, for excitons in QWs there is a translational invariance only in the two-dimensional $(x,y)$ QW plane. We introduce relative and center-of-mass coordinates (with $m_{\mathrm{h}}=m_{0}/\gamma_{1}$)
$x$ $=x_{\mathrm{e}}-x_{\mathrm{h}}\;,\;y=y_{\mathrm{e}}-y_{\mathrm{h}}\;,$ (7a)
$X$ $=\frac{m_{\mathrm{e}}x_{\mathrm{e}}+m_{\mathrm{h}}x_{\mathrm{h}}}{m_{\mathrm{e}}+m_{\mathrm{h}}}\;,\;Y=\frac{m_{\mathrm{e}}y_{\mathrm{e}}+m_{\mathrm{h}}y_{\mathrm{h}}}{m_{\mathrm{e}}+m_{\mathrm{h}}}$ (7b)
in this plane. Using relative and center-of-mass momenta, $\bm{p}_{(2D)}=(p_{1},p_{2})$ and $\bm{P}_{(2D)}=(P_{1},P_{2})$, respectively, and additionally setting the conserved center-of-mass momenta to zero, $P_{1}=P_{2}=0$, we obtain
$H(x,y,z_{\mathrm{e}},z_{\mathrm{h}},\bm{I},\bm{S}_{\text{h}})=E_{\text{g}}+H_{\text{SO}}-\frac{e^{2}}{4\pi\varepsilon_{0}\varepsilon}\frac{1}{\sqrt{x^{2}+y^{2}+(z_{\text{e}}-z_{\text{h}})^{2}}}+V_{\text{e}}(z_{\text{e}})+V_{\text{h}}(z_{\text{h}})+\frac{p^{2}_{\text{ez}}}{2m_{\text{e}}}$
$+\left[\frac{\gamma^{\prime}_{1}}{2m_{0}}+\frac{2\gamma_{2}}{m_{0}}+\frac{\eta_{1}+2\eta_{2}}{\hbar^{2}m_{0}}(\bm{I}\cdot\bm{S_{\text{h}}})\right]\bm{p}_{(2D)}^{2}+\left[\frac{\gamma_{1}}{2m_{0}}+\frac{2\gamma_{2}}{m_{0}}+\frac{\eta_{1}+2\eta_{2}}{\hbar^{2}m_{0}}(\bm{I}\cdot\bm{S_{\text{h}}})-\frac{3\gamma_{2}}{\hbar^{2}m_{0}}I_{3}^{2}-\frac{6\eta_{2}}{\hbar^{2}m_{0}}I_{3}S_{\text{h3}}\right]p_{\text{hz}}^{2}$
$-\frac{3\gamma_{2}}{\hbar^{2}m_{0}}\left(p_{1}^{2}I_{1}^{2}+p_{2}^{2}I_{2}^{2}\right)-\frac{6\gamma_{3}}{\hbar^{2}m_{0}}\left[\{p_{1},p_{2}\}\{I_{1},I_{2}\}\right]+\frac{6\gamma_{3}}{\hbar^{2}m_{0}}\left[\{p_{2},p_{\text{hz}}\}\{I_{2},I_{3}\}+\{p_{\text{hz}},p_{1}\}\{I_{3},I_{1}\}\right]$
$-\frac{6\eta_{2}}{\hbar^{2}m_{0}}(p_{1}^{2}I_{1}S_{\text{h1}}+p_{2}^{2}I_{2}S_{\text{h2}})-\frac{6\eta_{3}}{\hbar^{2}m_{0}}\left[\{p_{1},p_{2}\}(I_{1}S_{\text{h2}}+I_{2}S_{\text{h1}})\right]$
$+\frac{6\eta_{3}}{\hbar^{2}m_{0}}\left[\{p_{2},p_{\text{hz}}\}(I_{2}S_{\text{h3}}+I_{3}S_{\text{h2}})+\{p_{\text{hz}},p_{1}\}(I_{3}S_{\text{h1}}+I_{1}S_{\text{h3}})\right]$ (8)
with $\gamma^{\prime}_{1}=\gamma_{1}+m_{0}/m_{\mathrm{e}}$. When comparing the Hamiltonian (8) with the two-band model for excitons in a QW [36, 28, 29] it is important to note that not only the dimension of the Hamiltonian is increased by the additional spin degrees of freedom, i.e., the quasispin $\bm{I}$ and hole spin $\bm{S}_{\mathrm{h}}$ describing the valence band, but also the coordinate space is increased from three degrees of freedom $(\rho,z_{\mathrm{e}},z_{\mathrm{h}})$ in the two-band model to four degrees of freedom $(x,y,z_{\mathrm{e}},z_{\mathrm{h}})$ when considering the complete structure of the valence band. This is a consequence of the reduced symmetry of the Hamiltonian (8) as discussed in Sec. II.2.
The Hamiltonian in the form of Eq. (8) with Cartesian coordinates is not well suited for the computation of matrix elements in the basis introduced in Sec. II.4, which takes advantage of the polar coordinates $(\rho,\phi)$ in the QW plane and contains angular momentum functions $\exp(im\phi)$ and spin states $|J,m_{J}\rangle$ for the coupled quasispin and hole spin $\boldsymbol{J}=\boldsymbol{I}+\boldsymbol{S}_{\mathrm{h}}$. For this basis, it is more convenient to introduce creation and annihilation operators
$I_{\pm}$ $=I_{1}\pm iI_{2}\,,\quad S_{\mathrm{h}\pm}=S_{\mathrm{h}1}\pm iS_{\mathrm{h}2}\,,\quad p_{\pm}=p_{1}\pm ip_{2}\,$, (9)
where the operators $I_{\pm}$ and $S_{\mathrm{h}\pm}$ couple states with $\Delta m_{J}=\pm 1$ and the operators $p_{\pm}$ couple states with $\Delta m=\pm 1$. Representing the Hamiltonian (8) in terms of these operators and sorting terms with respect to their action on the angular momentum and spin part of the basis states, the complete Hamiltonian can finally be written as
$H=H_{\mathrm{diag}}+H_{\Delta m=0}+H_{\Delta m=1}+H_{\Delta m=2}+H_{\Delta m_{F}=4}$ (10)
with
$H_{\mathrm{diag}}=E_{\mathrm{g}}+H_{\mathrm{SO}}+V_{\mathrm{C}}+\frac{1}{2m_{\mathrm{e}}}p_{\mathrm{e}z}^{2}+\frac{\gamma_{1}^{\prime}}{2m_{0}}p_{+}p_{-}+\frac{\gamma_{1}}{2m_{0}}p_{\mathrm{h}z}^{2}+\left[\frac{2\gamma_{2}}{m_{0}}+\frac{\eta_{1}+2\eta_{2}}{\hbar^{2}m_{0}}(\boldsymbol{I}\cdot\boldsymbol{S}_{\mathrm{h}})\right]\left(p_{+}p_{-}+p_{\mathrm{h}z}^{2}\right)\,$, (11a)
$H_{\Delta m=0}=\left[-\frac{3\gamma_{2}}{2\hbar^{2}m_{0}}\{I_{+},I_{-}\}-\frac{3\eta_{2}}{2\hbar^{2}m_{0}}(I_{+}S_{\mathrm{h}-}+I_{-}S_{\mathrm{h}+})\right]p_{+}p_{-}+\left[-\frac{3\gamma_{2}}{\hbar^{2}m_{0}}I_{3}^{2}-\frac{6\eta_{2}}{\hbar^{2}m_{0}}I_{3}S_{\mathrm{h}3}\right]p_{\mathrm{h}z}^{2}\,$, (11b)
$H_{\Delta m=1}=\frac{3\gamma_{3}}{\hbar^{2}m_{0}}\left(\{I_{3},I_{+}\}p_{\mathrm{h}z}p_{-}+\mathrm{h.c.}\right)+\frac{3\eta_{3}}{\hbar^{2}m_{0}}\left[\left(I_{3}S_{\mathrm{h}+}+I_{+}S_{\mathrm{h}3}\right)p_{\mathrm{h}z}p_{-}+\mathrm{h.c.}\right]\,$, (11c)
$H_{\Delta m=2}=-\frac{3(\gamma_{2}+\gamma_{3})}{8\hbar^{2}m_{0}}\left(I_{+}^{2}p_{-}^{2}+\mathrm{h.c.}\right)-\frac{3(\eta_{2}+\eta_{3})}{4\hbar^{2}m_{0}}\left(I_{+}S_{\mathrm{h}+}p_{-}^{2}+\mathrm{h.c.}\right)\,$, (11d)
$H_{\Delta m_{F}=4}=-\frac{3(\gamma_{2}-\gamma_{3})}{8\hbar^{2}m_{0}}\left(I_{-}^{2}p_{-}^{2}+\mathrm{h.c.}\right)-\frac{3(\eta_{2}-\eta_{3})}{4\hbar^{2}m_{0}}\left(I_{-}S_{\mathrm{h}-}p_{-}^{2}+\mathrm{h.c.}\right)\,.$ (11e)
Here, $V_{\mathrm{C}}$ is the Coulomb potential and h.c. denotes the hermitian conjugate. The operator $H_{\mathrm{diag}}$ is diagonal with respect to the angular momentum and spin quantum numbers $m$, $J$, and $m_{J}=m_{I}+m_{S_{\mathrm{h}}}$. For an appropriately chosen effective hole mass this part of the Hamiltonian describes the exciton within a hydrogenlike two-band model as discussed in Refs. [36, 28, 29]. The operator $H_{\Delta m=0}$ couples basis states with quantum numbers $J=1/2$ and $J=3/2$ but the same angular momentum quantum number $m$, i.e., this term causes a coupling between the yellow and green exciton series without breaking the continuous rotational symmetry around the $z$ axis. The operators $H_{\Delta m=1}$ and $H_{\Delta m=2}$ couple states with different angular momentum quantum numbers $m$, and thus $m$ is not a good quantum number when considering the impact of the valence band. Finally, the operator $H_{\Delta m_{F}=4}$ couples states with $\Delta m_{F}=\pm 4$, where $m_{F}=m+m_{J}$ is the quantum number for the $z$ component of the coupled spin and angular momenta $\boldsymbol{F}=\boldsymbol{L}+\boldsymbol{J}=\boldsymbol{L}+\boldsymbol{I}+\boldsymbol{S}_{\mathrm{h}}$. Note that when using basis states with given quantum numbers $m_{F}$ the matrix representation of the Hamiltonian (10) has a block structure with four blocks, which can be diagonalized separately. This block structure is directly related to the symmetry properties of the Hamiltonian (10) as shown below in Sec. II.2. Note also that $H_{\Delta m_{F}=4}$ would vanish in case of equal Luttinger parameters $\gamma_{2}=\gamma_{3}$ and $\eta_{2}=\eta_{3}$, i.e., $m_{F}$ would be an exact quantum number in that case.
II.2 Symmetry properties
In comparison to the hydrogenlike model of excitons in a QW [29], the inclusion of the complex valence band leads to a further reduction of the symmetry. We are considering the situation in which the [001] axis of the crystal is perpendicular to the QW plane. In this case, the symmetry of the system is reduced from $D_{\infty\mathrm{h}}$, and the complete Hamiltonian (10) is now only invariant under the symmetry operations of the $D_{4\mathrm{h}}$ symmetry group [46], containing rotations $C_{4}$ and $C_{2}$, inversion $I$, improper rotations $S_{4}$, and reflections $\sigma_{\mathrm{h}}$, $\sigma_{\mathrm{v}}$, and $\sigma_{\mathrm{d}}$, as illustrated in Fig. 3. With $m_{J}=m_{I}+m_{S_{\mathrm{h}}}$ and $m_{F}=m+m_{J}$, the $C_{4}$ symmetry implies that the Hamiltonian only couples basis states with $\Delta m_{F}=0,\pm 4$. Note that this symmetry-derived constraint on matrix elements only applies to the total angular momentum, i.e., individual angular momenta like $m$ or $m_{J}$ are not restricted in this way in their possible couplings, as can be seen by the terms in the Hamiltonian (10).
Figure 3: Illustration of symmetry operations of the group $D_{4\mathrm{h}}$. $C_{2}$ and $C_{4}$ denote two- and fourfold rotations, respectively, $\sigma_{h}$, $\sigma_{v}$, and $\sigma_{d}$ reflections with respect to different kinds of symmetry planes. In addition to the depicted operations, there is also the inversion II and the improper rotation $S_{4}=\sigma_{h}C_{4}$.
Due to the reduced symmetry of the system, the eigenstates can no longer be classified according to their angular momentum quantum numbers, but instead by the irreducible representations of the symmetry group $D_{4\mathrm{h}}$, or rather, as there are also spin degrees of freedom, the corresponding double group to $D_{4\mathrm{h}}$. Without consideration of the electron spin $\boldsymbol{S}_{\mathrm{e}}$, the total angular momentum $\boldsymbol{F}$ in the system is half-integer valued, and only the irreducible representations $\Gamma_{6}^{+}$, $\Gamma_{6}^{-}$, $\Gamma_{7}^{+}$, and $\Gamma_{7}^{-}$ are possible for the eigenstates [46], where the superscript ++ or - indicates the total parity of the corresponding eigenstates under inversion $I$. The eigenstates are superpositions of states with $m_{F}$ differing by multiples of 4, and thus can be assigned to one of the following groups
$m_{F}$ $=\left\{\begin{array}[]{l}\dots,-\frac{13}{2},-\frac{5}{2},\frac{3}{2},\frac{11}{2},\dots\\[4.30554pt] \dots,-\frac{11}{2},-\frac{3}{2},\frac{5}{2},\frac{13}{2},\dots\end{array}\right$. (12c)
$m_{F}$ $=\left\{\begin{array}[]{l}\dots,-\frac{15}{2},-\frac{7}{2},\frac{1}{2},\frac{9}{2},\dots\\[4.30554pt] \dots,-\frac{9}{2},-\frac{1}{2},\frac{7}{2},\frac{15}{2},\dots\end{array}\right$. (12f)
Here, blocks related by a sign change of $m_{F}$ are degenerate due to time-reversal symmetry. Equations (12c) and (12f) describe twofold degenerate states in subspaces $\Gamma_{6}^{\pm}$ and $\Gamma_{7}^{\pm}$, respectively, as can be seen as follows.
To understand how these blocks relate to the irreducible representations $\Gamma_{6/7}^{\pm}$, we can look at a state with total angular momentum $F=1/2$. We need to take into account that the eigenstates of the quasispin $I$ do not behave exactly like eigenstates of a standard angular momentum. In $D_{4\mathrm{h}}$, the eigenstates of the quasispin transform like the sum of irreducible representations $\Gamma_{4}^{+}+\Gamma_{5}^{+}=(\Gamma_{2}^{+}+\Gamma_{5}^{+})\times\Gamma_{3}^{+}$, while the eigenstates of a standard spin of 1 behave like $\Gamma_{2}^{+}+\Gamma_{5}^{+}$. Hence, one obtains the total symmetry by the usual addition of angular momenta, followed by a multiplication with $\Gamma_{3}^{+}$. With this prescription in mind, we can use that an angular momentum of $1/2$ transforms according to $\Gamma_{6}^{\pm}$ in $D_{4\mathrm{h}}$ [46]. Multiplying by $\Gamma_{3}^{+}$, we find that a state with total angular momentum $F=1/2$, and thus states with $m_{F}=1/2$ and $m_{F}=-1/2$ transform according to $\Gamma_{7}^{\pm}$.
For the computation of oscillator strengths in Sec. II.3, the electron spin $\boldsymbol{S}_{\mathrm{e}}$ must also be taken into account. The total symmetries of the system including the electron spin are discussed below in Sec. II.3.1.
II.3 Oscillator strengths
For the computation of oscillator strengths we generalize the theory developed in Ref. [29] within a two-band model. We consider excitons in the many-particle framework. The ground state $|\Phi_{0}\rangle$ of the crystal is the state where all electrons of the crystal are in the valence bands, while all conduction bands are empty. An exciton is created when one of these electrons is lifted from the valence bands into a conduction band. The resulting state can be written as
$|\psi^{\sigma\tau}_{vc,\nu}\rangle=\sum_{\boldsymbol{q}_{c},\boldsymbol{q}_{v}}f^{\tau,\sigma}_{vc,\nu}(\boldsymbol{q}_{c},\boldsymbol{q}_{v})c^{\dagger}_{c,\boldsymbol{q}_{c},\tau}c_{v,\boldsymbol{q}_{v},\sigma}|\Phi_{0}\rangle\,.$ (13)
Here, $c^{\dagger}_{c,\boldsymbol{q}_{c}}$ denotes the creation operator for an electron with crystal momentum $\boldsymbol{q}_{c}$ and electron spin $\tau$ in the conduction band $c$, and $c_{v,\boldsymbol{q}_{v}}$ is the analogous annihilation operator in the valence band $v$, which creates a hole with crystal momentum $\boldsymbol{q}_{v}$ and hole spin $\sigma$. The coefficient $f^{\tau,\sigma}_{vc,\nu}(\boldsymbol{q}_{c},\boldsymbol{q}_{v})$ in the superposition is the envelope function, which is obtained as the solution of the exciton Schrödinger equation with Hamiltonian (10) transformed into crystal-momentum space. When considering the complex valence band structure, $c$ and $v$ are not good quantum numbers, and the correct states are obtained by forming superpositions. The subscript $\nu$ denotes further quantum numbers or indices associated with the state.
[…]
We are specifically interested in the oscillator strengths for circularly polarized light. We thus need to take suitable linear combinations of the transition matrix elements in Eqs. (14)
[…]
Other polarizations can be evaluated analogously.
Using these relations and the Clebsch-Gordan coefficients, the spin and quasispin bra vectors can be rewritten as
[…]
with the coefficients
[…]
$=\;$ $\sum_{m}\int_{-\frac{L}{2}}^{\frac{L}{2}}\psi_{m,J,m_{J},m_{S_{\mathrm{e}}}}^{\nu\boldsymbol{K}_{\mathrm{2D}}}(\rho,z,Z)\frac{e^{im\phi}}{\sqrt{2\pi L}}e^{-iK_{z}Z}\mathrm{d}Z.$ (19)
The photon momenta for dipole transitions to yellow excitons ($\lambda\approx 580\,$nm) are negligibly small. We can thus set $K_{z}$ and $\boldsymbol{K}_{\mathrm{2D}}$ to zero and ignore it in the following. Common prefactors can also be dropped in the calculation of relative oscillator strengths. From the condition of continuity at the origin, we can deduce for the derivatives that [29]
$\frac{\partial}{\partial x}\left(\psi_{m,J,m_{J},m_{S_{\mathrm{e}}}}^{\nu}(\rho,z,Z)e^{im\phi}\right)\Bigr\rvert_{\rho=z=0}$
[…]
$\frac{\partial}{\partial y}\left(\psi_{m,J,m_{J},m_{S_{\mathrm{e}}}}^{\nu}(\rho,z,Z)e^{im\phi}\right)\Bigr\rvert_{\rho=z=0}$
[…]
$\frac{\partial}{\partial z}\left(\psi_{m,J,m_{J},m_{S_{\mathrm{e}}}}^{\nu}(\rho,z,Z)e^{im\phi}\right)\Bigr\rvert_{\rho=z=0}$
$=\;$ $\delta_{m,0}\partial_{z}\psi_{m,J,m_{J},m_{S_{\mathrm{e}}}}^{\nu}(\rho,z,Z)\bigr\rvert_{\rho=z=0}\,.$ (20c)
Combining Eqs. (16), (19), and (20), the desired oscillator strengths can be evaluated as
$f^{\mathrm{rel}}_{\pm}=$ $\bigg|\int_{-\frac{L}{2}}^{\frac{L}{2}}\left[\mathcal{M}_{\rho}^{\pm}(Z)+\mathcal{M}_{z}^{\pm}(Z)\right]\mathrm{d}Z\bigg|^{2}$ (21)
with the transition amplitude densities
$\mathcal{M}^{\pm}_{\rho}(Z)$ $=\sqrt{2}\sum_{J,m_{J}}c^{m_{I}=0,\ast}_{J,m_{J},m_{S_{\mathrm{e}}}}\partial_{\rho}\psi_{\pm 1,J,m_{J},m_{S_{\mathrm{e}}}}^{\nu}(\rho,z,Z)\Bigr\rvert_{\rho=z=0}\,$, (22a)
$\mathcal{M}^{\pm}_{z}(Z)$ $=\sum_{J,m_{J}}c^{m_{I}=\pm 1,\ast}_{J,m_{J},m_{S_{\mathrm{e}}}}\partial_{z}\psi_{0,J,m_{J},m_{S_{\mathrm{e}}}}^{\nu}(\rho,z,Z)\Bigr\rvert_{\rho=z=0}\,$, (22b)
and the complex conjugate of the coefficients given in Eq. (18). Details on the numerical evaluation of Eqs. (21) and (22) are presented in Sec. II.4.
Note that in case of degenerate eigenenergies, the oscillator strengths calculated with Eq. (21) for the individual degenerate states must be added.
II.3.1 Selection rules for dipole transitions
As the dipole operator has odd parity, and parity is a good quantum number in the system, we can deduce that only states with odd total parity belonging to the irreducible representations $\Gamma_{6}^{-}$ and $\Gamma_{7}^{-}$ can be optically excited in the setup shown in Fig. 1. Correspondingly, the integrals in Eq. (21) must vanish for states with even total parity.
Precise selection rules can be deduced by application of group theory [46] when considering the symmetry of the dipole operator and the respective symmetries of the states. In the context of oscillator strengths, we add the electron spin symmetry $\Gamma_{6}^{+}$, leading to the total symmetries that decompose into a sum of subspaces as
$\Gamma_{6}^{\pm}\times\Gamma_{6}^{+}$ $=\Gamma_{1}^{\pm}+\Gamma_{2}^{\pm}+\Gamma_{5}^{\pm}\,$, (23a)
$\Gamma_{7}^{\pm}\times\Gamma_{6}^{+}$ $=\Gamma_{3}^{\pm}+\Gamma_{4}^{\pm}+\Gamma_{5}^{\pm}\,.$ (23b)
The dipole operator $D^{-}_{l=1}$ on the other hand decomposes into the sum $\Gamma_{2}^{-}+\Gamma_{5}^{-}$ of irreducible representations. Here, $\Gamma_{2}^{-}$ corresponds to linear polarization along the $z$ direction, whereas the two-dimensional $\Gamma_{5}^{-}$ corresponds to polarizations in the QW plane, including the circular polarizations we consider here. From this, it follows that light circularly polarized in the $xy$ plane can create excitons in the twofold degenerate $\Gamma_{5}^{-}$ spaces. As $m_{S_{\mathrm{e}}}$ is a good quantum number, the entire negative-parity subspaces in Eqs. (23) are degenerate. As they both fully contain $\Gamma_{5}^{-}$ spaces, it follows that they can both be excited by left- and right-handed circularly polarized light. In fact, from time-reversal symmetry of the system, it follows that both polarizations lead to identical spectra.
Note that a more restricted selection rule is valid for $z$-polarized light. In that case, only $\Gamma_{6}^{-}$ states can be excited. However, this would require light oriented in the QW plane, and thus a different experimental setup than shown in Fig. 1.
II.4 Numerical algorithm
To compute the exciton eigenenergies and eigenfunctions of the Hamiltonian (10), we extend the basis used in the hydrogenlike model [36, 28] for the wave functions in the three-dimensional coordinate space $(\rho,z_{\mathrm{e}},z_{\mathrm{h}})$ by the additional angular momentum and spin degrees of freedom. We use the ansatz
[…]
with the basis states
$|m,J,m_{J}\rangle=e^{im\phi}\,|J,m_{J}\rangle$ (25)
as the product of the eigenstates of the angular momentum operator $L_{z}$ and the eigenstates of the coupled quasispin and hole spin $\bm{J}=\bm{I}+\bm{S}_{\mathrm{h}}$. The wave functions $\psi_{m,J,m_{J}}(\rho,z_{\mathrm{e}},z_{\mathrm{h}})$ in Eq. (24) are expanded over a basis of B-splines of order $n$ as
$\psi_{m,J,m_{J}}(\rho,z_{\mathrm{e}},z_{\mathrm{h}})$
$=$ $\frac{1}{\sqrt{\rho}}\sum_{i,j,k}c^{J,m_{J},m}_{ijk}\,B_{i}^{n}(\rho)\,B_{j}^{n}(z_{\mathrm{e}})\,B_{k}^{n}(z_{\mathrm{h}})$ (26)
similar as in Refs. [36, 28]. For convenience, the definition and some properties of B-splines are collected in Appendix A. Note that we do not expand the $\phi$ dependence in Eq. (24) using B-splines but rather use the angular momentum states $\exp(im\phi)$ as basis states.
The expansions (24) and (26) lead to a generalized eigenvalue problem
$\sum_{\beta}H_{\alpha\beta}c_{\beta}=E\sum_{\beta}O_{\alpha\beta}c_{\beta}$ (27)
for the energy eigenvalues of the Hamiltonian (10) and the expansion coefficients $c^{J,m_{J},m}_{ijk}$ of the corresponding eigenfunctions. Here, the indices $\alpha$ and $\beta$ count both the elements of the total basis set composed of B-splines, angular momentum, and spin states in Eq. (24) and the expansion coefficients $c^{J,m_{J},m}_{ijk}$ when arranged as one-dimensional vectors.
The elements of the matrices in Eq. (27) are computed as follows. The matrix elements of all terms of the Hamiltonian (10), except for the Coulomb potential $V_{C}$, can be factorized into products of one-dimensional integrals of B-splines or their derivatives along coordinates $\rho$, $z_{\mathrm{e}}$, and $z_{\mathrm{h}}$, as well as analytically pre-calculated matrix elements of spin operators. Details of the computation of matrix elements of the momentum operators and the spin operators are given in Appendices B and C, respectively. The matrix elements of the Coulomb potential $V_{C}$ are diagonal with respect to the angular momentum and spin basis. The non-trivial three-dimensional integrals in the $(\rho,z_{\mathrm{e}},z_{\mathrm{h}})$ coordinate space are computed, using the B-spline basis, as explained in Ref. [36].
Table 2: Selected angular momentum and spin basis states $|m,J,m_{J}\rangle$ with quantum numbers $m_{F}=-7/2$, $1/2$, and $9/2$ belonging to the irreducible subspaces $\Gamma_{7}^{\pm}$ and used for numerical calculations.
$m_{F}=-\frac{7}{2}$ $m_{F}=\frac{1}{2}$ $m_{F}=\frac{9}{2}$
$|{-5},\frac{3}{2},+\frac{3}{2}\rangle$ $|{-1},\frac{3}{2},+\frac{3}{2}\rangle$ $|3,\frac{3}{2},+\frac{3}{2}\rangle$
$|{-4},\frac{3}{2},+\frac{1}{2}\rangle$ $|{+0},\frac{3}{2},+\frac{1}{2}\rangle$ $|4,\frac{3}{2},+\frac{1}{2}\rangle$
$|{-4},\frac{1}{2},+\frac{1}{2}\rangle$ $|{+0},\frac{1}{2},+\frac{1}{2}\rangle$ $|4,\frac{1}{2},+\frac{1}{2}\rangle$
$|{-3},\frac{1}{2},-\frac{1}{2}\rangle$ $|{+1},\frac{1}{2},-\frac{1}{2}\rangle$ $|5,\frac{1}{2},-\frac{1}{2}\rangle$
$|{-3},\frac{3}{2},-\frac{1}{2}\rangle$ $|{+1},\frac{3}{2},-\frac{1}{2}\rangle$ $|5,\frac{3}{2},-\frac{1}{2}\rangle$
$|{-2},\frac{3}{2},-\frac{3}{2}\rangle$ $|{+2},\frac{3}{2},-\frac{3}{2}\rangle$ $|6,\frac{3}{2},-\frac{3}{2}\rangle$
The generalized eigenvalue problem (27) is solved using ARPACK routines optimized for band matrices [51]. The band and even sparse structure of the resulting Hamiltonian matrix $H_{\alpha\beta}$ and overlap matrix $O_{\alpha\beta}$ arises naturally from the compact support of the B-spline functions. We use 15 B-spline functions of order $n=5$ for both the $z_{\mathrm{e}}$ and $z_{\mathrm{h}}$ directions, defined on equidistant grids, and 26 B-spline functions for the $\rho$ direction. For the $\rho$ coordinate, the grid nodes are distributed non-equidistantly with a decreasing density over the interval $0,200\,\mathrm{nm}]$, providing a finer resolution near the origin where both the potential and the wave function vary most rapidly, see Appendix A for more details. Because only basis states with $\Delta m_{F}=0,\pm 4$ are coupled, it is sufficient to use one of the sets in Eq. (12) for the numerical diagonalization of the Hamiltonian.
We are mainly interested in the optically active states with dominant $m=0$ and $m=1$ contributions. Convergence of these states is achieved when restricting the spin and angular momentum basis to the subset with $m_{F}=-7/2$, $1/2$, and $9/2$ given in Table 2 in the irreducible subspace $\Gamma_{7}^{\pm}$ and to the subset with $m_{F}=-5/2$, $3/2$, and $11/2$ in the subspace $\Gamma_{6}^{\pm}$. The total number $N$ of basis states and thus the dimension of the generalized eigenvalue problem (27) is given by the product of the number of B-splines for the $\rho$, $z_{\mathrm{e}}$, and $z_{\mathrm{h}}$ coordinates and the number of angular momentum and spin states $|m,J,m_{J}\rangle$, and was up to $N=26\times 15\times 15\times 18=105\,300$ in our calculations.
Figure 4: Exciton spectra in a cuprous oxide QW. In (a) the QW width is increased from $L=0$ to 1010\,nm for excitons described by the hydrogenlike part of the Hamiltonian given by the first two terms in Eq. (30). In (b)-(e) the non-diagonal coupling terms of the are successively switched on using the parameters $\lambda_{0,1,2,3}\in[0,1]$ in the Hamiltonian (30) (see text). The photoabsorption spectrum for transitions with circularly polarized light and considering the complete structure of the valence band is given in (f). Some states are labeled by approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$. The colors of symbols or lines indicate the expectation value $\langle m\rangle$ of the angular momentum of the exciton state.
To compute oscillator strengths we use the eigenstates, obtained by solving the generalized eigenvalue problem. The solution yields the wave function for each state in the coordinates $\rho$, $z_{\mathrm{e}}$, and $z_{\mathrm{h}}$. For evaluating Eqs. (22), the wave functions have to be expressed in relative coordinates in the $z$ direction. As we only need the wave function at $z=0$, i.e., at $z_{\mathrm{e}}=z_{\mathrm{h}}=Z$, the transformation is straightforward and given by
$\psi(\rho,z\mathord{=}0,Z)=\psi(\rho,z_{\mathrm{e}}\mathord{=}Z,z_{\mathrm{h}}\mathord{=}Z).$ (28)
The derivative with respect to $z$ is calculated analytically by differentiating the B-spline basis functions according to Eq. (31). It acts on a product of basis states as
$\partial_{z}B_{j}^{n}(z_{\mathrm{e}})\,B_{k}^{n}(z_{\mathrm{h}})=\partial_{z_{\mathrm{e}}}B_{j}^{n}(z_{\mathrm{e}})\,B_{k}^{n}(z_{\mathrm{h}})-B_{j}^{n}(z_{\mathrm{e}})\,\partial_{z_{\mathrm{h}}}B_{k}^{n}(z_{\mathrm{h}}).$ (29)
By contrast, the derivative with respect to $\rho$ at $\rho=0$ is computed using the approximation $\partial_{\rho}\psi(\rho,z\mathord{=}0,Z)\approx\frac{1}{\rho}\psi(\rho,z\mathord{=}0,Z)$, which is extrapolated, analogously to Ref. [29], to $\rho=0$. To compute the integral in Eq. (21), the integrands are evaluated at 100 positions on a equidistant grid, and the integral is solved numerically using numpy.trapz [52]. Note that the most time consuming part is the diagonalization of the generalized eigenvalue problem (27).
III Results and discussion
The complex structure of the cuprous oxide valence bands has a significant effect on excitons in the QW. To observe the shifts of energy levels due to the impact of the different terms in the Hamiltonian (10), we apply the following strategy. We start by taking into account the diagonal part of the Hamiltonian, $H_{\mathrm{diag}}$, for the hydrogenlike two-band model of excitons in a QW, and then switch on, term by term, the parts in the Hamiltonian, which are non-diagonal with respect to the angular momentum and spin quantum numbers of the basis. To this aim, we modify the Hamiltonian (10) to the form
$H$ $=H_{\mathrm{diag}}+H_{\Delta m=0}\delta_{J^{\prime}J}+{\lambda_{0}}H_{\Delta m=0}(1-\delta_{J^{\prime}J})$
$+{\lambda_{1}}H_{\Delta m=1}+{\lambda_{2}}H_{\Delta m=2}+{\lambda_{3}}H_{\Delta m_{F}=4}$ (30)
by introducing the auxiliary parameters $\lambda_{0,1,2,3}\in[0,1]$, which can be used to successively switch on the individual non-diagonal terms. Note that $H_{\Delta m=0}$ in Eq. (11b) has both, nonzero diagonal and non-diagonal terms in the angular momentum and spin basis $|m,J,m_{J}\rangle$. In Eq. (30), the nonzero diagonal terms of $H_{\Delta m=0}$ are considered in the hydrogenlike model, described by the first and second term, and the non-diagonal terms, which couple $J=1/2$ and $J=3/2$ states of the yellow and green exciton series, respectively, are switched on by increasing $\lambda_{0}$ from 0 to $1$. For an energy range of the lowest, yellow exciton series a sequence of spectra obtained by diagonalization of the Hamiltonian (30) is presented in Fig. 4. The figure allows for a detailed discussion of the impact of the valence band on excitons in a cuprous oxide QW of L=10L=10\,nm width.
We go through Fig. 4 from left to right. Panel (a) shows spectra of the hydrogenlike two-band model for excitons in a QW of width $L$ in the range from $L=0$ to L=10L=10\,nm. The energies are given with respect to the first scattering threshold $E_{1,1}(L)$, given by the sum of the lowest quantum-confinement energies, see Eq. (6). Note that the QW width $L=0$ refers to the analytically solvable two-dimensional Coulomb problem [30, 31], where states with the same principal quantum number $n$ but different angular momentum quantum numbers $m$ are degenerate, as can be clearly seen in Fig. 4. The degeneracy is lifted, when $L$ is increased, in the transition region from strong to weak confinement [28]. Some of the states in panel (a) are marked by quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$, with $m$ being an exact quantum number due to the rotational symmetry of the hydrogenlike problem. The colors of the states indicate the expectation values $\langle m\rangle$ of the angular momentum operator throughout Fig. 4.
In panels (b)–(f) the QW width is fixed to L=10L=10\,nm to study the impact of the complex valence band on the excitons at such a QW width. In panel (b) the non-diagonal terms of $H_{\Delta m=0}$, coupling the yellow ($J=1/2$) and green ($J=3/2$) excitons, are switched on by increasing $\lambda_{0}$ from 0 to $1$. The coupling between yellow and green excitons has a strong impact on states with angular quantum number $m=0$, and smaller effect on states with $m\neq 0$, the energies of those are moderately lowered. Note that for the states in this panel the rotational symmetry around the $z$ axis is still valid, and thus $m$ and $m_{J}$ are exact quantum numbers. As the eigenenergies do not depend on the sign of these quantum numbers, states with $m\neq 0$ are fourfold degenerate (note that $m_{F}$ is half-integer).
In panels (c) and (d) of Fig. 4 the non-diagonal terms $H_{\Delta m=1}$ and $H_{\Delta m=2}$ in the Hamiltonian (30) are successively switched on. These coupling terms break the rotational symmetry around the $z$ axis in a way that $m$ is not a good quantum number any longer, but the quantum number $m_{F}=m+m_{J}$ for the $z$ component of the total angular momentum $\boldsymbol{F}=\boldsymbol{L}+\boldsymbol{J}$ still is. As an important consequence, the degeneracy of, e.g., some states with $m_{F}=1/2$ and $m_{F}=3/2$, marked by solid and open symbols, respectively, is reduced in panels (c) and (d) of Fig. 4 from fourfold to twofold degeneracy.
Finally, in panel (e) the non-diagonal terms $H_{\Delta m_{F}=4}$ in the Hamiltonian (30), which couple basis states with $\Delta m=\pm 2$ and $\Delta m_{F}=\pm 4$ are switched on. In this panel, $m_{F}$ is not a good quantum number any longer, and states can only be classified as belonging to one of the four irreducible subspaces $\Gamma_{6}^{\pm}$ and $\Gamma_{7}^{\pm}$, as discussed in Sec. II.2. The eigenenergies are shifted due to the coupling between different $m_{F}$ states, however, there is no further lifting of degeneracies, i.e., all states remain twofold degenerate. The rightmost spectrum in Fig. 4(e) is the one with $\lambda_{0,1,2,3}=1$, where the full impact of the valence band as described by the Hamiltonian (10) is taken into account. The oscillator strengths corresponding to this spectrum are shown in Fig. 4(f). Note that the spectrum significantly differs from that of the hydrogenlike model shown at QW width L=10L=10\,nm in Fig. 4(a). This clearly illustrates that a precise description of exciton states in QWs requires taking into account the complex structure of the valence band to accurately determine all splittings of energy levels.
Figure 5: Photoabsorption spectrum of a cuprous oxide quantum well of width L=10L=10\,nm for transitions with left-handed circularly polarized light ($\hat{\bm{e}}_{x}+i\hat{\bm{e}}_{y}$). The relative oscillator strengths $f_{+}^{\mathrm{rel}}$ of parity allowed transitions to states in the subspaces $\Gamma_{6}^{-}$ and $\Gamma_{7}^{-}$ are represented by vertical light blue and green bars, respectively. Parity forbidden even states in the subspaces $\Gamma_{6}^{+}$ and $\Gamma_{7}^{+}$ are marked by dark blue and dark green arrows. The insets show the transition amplitude density $\mathcal{M}_{\rho}^{+}(Z)$ (orange lines) and $\mathcal{M}_{z}^{+}(Z)$ (red lines) given in Eqs. (22a) and (22b), respectively, together with their sums (black lines). Some states are marked by approximate quantum numbers $(n,m,N_{\mathrm{e}},N_{\mathrm{h}})$.
For a more detailed discussion, the photoabsorption spectrum of a cuprous oxide QW of width L=10L=10\,nm is shown in Fig. 5. Excitons are excited by a laser beam oriented perpendicular to the QW plane with left-handed ($\hat{\bm{e}}_{x}+i\hat{\bm{e}}_{y}$) circularly polarized light. Note that due to the selection rules discussed in Sec. II.3.1 only states in the irreducible subspaces $\Gamma_{6}^{-}$ and $\Gamma_{7}^{-}$ can be excited. Their relative oscillator strengths $f_{+}^{\mathrm{rel}}$ are computed by evaluation of Eq. (21) and indicated in Fig. 5 by vertical light green and light blue bars, respectively. Parity-forbidden even states in the subspaces $\Gamma_{6}^{+}$ and $\Gamma_{7}^{+}$ are marked by dark blue and dark green arrows. The insets show the transition amplitude density $\mathcal{M}_{\rho}^{+}(Z)$ (orange lines) and $\mathcal{M}_{z}^{+}(Z)$ (red lines) given in Eqs. (22a) and (22b), respectively. As can be clearly seen, they are even functions in $Z$ for the parity allowed and odd functions for the parity forbidden states. It should be noted that, contrary to the simple two-band model discussed in Ref. [29], both densities $\mathcal{M}_{\rho}^{+}(Z)$ and $\mathcal{M}_{z}^{+}(Z)$ are nonzero and constructively or destructively interfere to the sums shown as black lines in the insets of Fig. 5.
IV Conclusion and outlook
We have studied the impact of the valence band structure on the energies and level splittings of yellow excitons in cuprous oxide QWs. Based on the Luttinger-Kohn model, we derived the complete Hamiltonian, taking into account the full complex valence band structure. We calculated the exciton energies by expanding the wave function over a multidimensional basis of angular and spin eigenstates and the B-splines for the coordinate part of the wave function. This allowed us to observe the breaking of the rotational symmetry, leading to the quantum number $m$ no longer being a good quantum number. The step-by-step inclusion of the nondiagonal coupling terms of the complex valence band Hamiltonian clearly demonstrated shifts and splittings of the energy levels. We also computed the relative oscillator strengths for the excitonic transitions in QWs with circularly polarized light.
In this article, we have considered the complete valence band structure but have still used some approximations and simplifications, which shall be addressed in future work. In the Hamiltonian (8) we have used the same Luttinger parameters as in the bulk material [10]. This may be a good approximation for sufficiently broad QWs ($L\gtrsim 5$ nm), however, the Luttinger parameters probably need to be adjusted for narrow QWs based on density functional theory (DFT) calculations [53] of the electronic band structure of the thin films [54]. In our calculations we have also assumed the same dielectric constant $\varepsilon$ for both the crystal and the substrate. The dielectric contrast of a more realistic setup can be taken into account via the Rytova-Keldysh potential [42, 43] similar to what has been done for the hydrogenlike model of excitons in QWs [28]. Furthermore, we have restricted our computations to a configuration in which the QW is aligned perpendicular to the $[001]$ axis of the crystal, and excitons are invariant under symmetry transformations of the $D_{4\mathrm{h}}$ symmetry group. A more general orientation between the QW and the crystal can be considered, however, this will further reduce the symmetries of the problem. For the computation of spectra and oscillator strengths we have focused on bound states of the yellow exciton series. In future work, this can be generalized to yellow and green resonance states above threshold by application of the complex-coordinate-rotation method as in Ref. [36]. This will also enable the search for BICs [37]. The consideration of all these effects will certainly allow for detailed comparisons with future experimental spectra.
Appendix A B-splines
B-splines of higher orders are an effective tool for discretizing and solving partial differential equations. The unknown function is expanded over a basis of B-splines $B^{k}_{i}(x)$, $i=1,\ldots,\mathcal{N}$, which are piecewise polynomials of degree k1k-1. Given the predefined series of the service nodes $t_{i}$, each B-spline $B^{k}_{i}(x)$ of order $k$ is defined on the interval $x\in[t_{i},t_{i+k}]$. Values of the B-splines $B^{k}_{i}(x)$ and their derivatives at a given point $x$ can be calculated by recursion formulas [41]
$B^{k}_{i}(x)$ $=\frac{x-t_{i}}{t_{i+k-1}-t_{i}}B^{k-1}_{i}(x)+\frac{t_{i+k}-x}{t_{i+k}-t_{i+1}}B^{k-1}_{i+1}(x)\,$,
$\frac{dB^{k}_{i}(x)}{dx}$ $=\frac{k-1}{t_{i+k-1}-t_{i}}B^{k-1}_{i}(x)-\frac{k-1}{t_{i+k}-t_{i+1}}B^{k-1}_{i+1}(x)\,.$ (31)
The expression for the second derivative can be derived from the above two equations.
B-splines have some important characteristics. First, a B-spline of order $k$ on an equidistant grid approximates an analytical function with accuracy of $h^{k}$, where $h$ is the step size of the grid. Thus, higher-order B-splines give an accurate solution even for a relatively small number of nodes, which is particularly important for multi-dimensional problems. Second, one can choose nodes non-equidistantly and add service nodes at the boundaries in such a way as to have some B-splines equal to 1 at the boundaries, while all other B-splines are exactly zero there. Then, for example, zero boundary conditions can be easily implemented by removing the B-splines which are nonzero at the boundaries. Third, the B-splines are nonorthogonal functions, and thus the problem turns into a generalized eigenvalue problem. However, the B-spline functions have minimal support, i.e., each B-spline vanishes, $B^{k}_{i}(x)=0$, for $x\notin[t_{i},t_{i+k}]$, which significantly reduces the number of integrations to calculate matrix elements. This leads to a sparse structure of the matrices of the generalized eigenvalue problem (27).
Appendix B Matrix elements of momentum operators
In polar coordinates $(\rho,\phi)$ the creation and annihilation operators $p_{\pm}$ are given as
$p_{\pm}=\frac{\hbar}{i}\exp(\pm i\phi)\left[\partial_{\rho}\mp\frac{1}{\rho}\left(\frac{L_{z}}{\hbar}\pm\frac{1}{2}\right)\right]$ (32)
with $p_{+}=p_{-}^{\dagger}$. Using Eq. (32) and the commutator $L_{z}/\hbar,\exp(\pm i\phi)]=\pm 1$ the squared operators $p_{\pm}^{2}$ are obtained as
$p_{\pm}^{2}=$ $-\hbar^{2}\exp(\pm 2i\phi)\left[\partial_{\rho}^{2}\mp\frac{2}{\rho}\partial_{\rho}\left(\frac{L_{z}}{\hbar}\pm 1\right)\right$.
$+\left.\frac{1}{\rho^{2}}\left(\frac{L_{z}}{\hbar}\pm\frac{5}{2}\right)\left(\frac{L_{z}}{\hbar}\pm\frac{1}{2}\right)\right]\,.$ (33)
The operator $p_{+}p_{-}$ reads
$p_{+}p_{-}=-\hbar^{2}\partial_{\rho}^{2}+\frac{1}{\rho^{2}}\left(L_{z}^{2}-\frac{\hbar^{2}}{4}\right)\,$, (34)
and thus is diagonal with respect to the angular momentum functions $\exp(im\phi)$. For functions given as a product of a B-spline over the $\rho$ coordinate and an angular function $\exp(im\phi)$, it is now straightforward to evaluate matrix elements of the operators given above. The $\phi$ integrals are trivially solved analytically, yielding selection rules for non-zero matrix elements between states with quantum numbers $m$ and $m^{\prime}$. The B-spline functions as well as their derivatives are piecewise polynomials, and can be easily integrated. Note that due to the boundary condition that the B-splines vanish at $\rho=0$, the product of two B-splines can be divided by $\rho$ or $\rho^{2}$ without the occurrence of any singularity at the origin.
Using B-splines over the coordinates $z_{\mathrm{e}}$ and $z_{\mathrm{h}}$, matrix elements of the operators $p_{\mathrm{e}z}^{2}=-\hbar^{2}\partial_{\mathrm{e}z}^{2}$, $p_{\mathrm{h}z}=(\hbar/i)\partial_{\mathrm{h}z}$, and $p_{\mathrm{h}z}^{2}=-\hbar^{2}\partial_{\mathrm{h}z}^{2}$ are straightforwardly computed by integration of products of B-splines and their derivatives.
Appendix C Matrix elements of spin operators
The matrix elements of the quasispin and hole spin operators $\bm{I}$ and $\bm{S}_{\mathrm{h}}$ are computed in the six-dimensional basis $|J,m_{J}\rangle$ of the spin operator $\bm{J}=\bm{I}+\bm{S}_{\mathrm{h}}$, which can also be expressed, using Clebsch-Gordan coefficients, in the basis $|I,m_{I};S_{\mathrm{h}},m_{S_{\mathrm{h}}}\rangle$, i.e.
[…]
With these basis states the matrix representations of the spin operators are obtained as
$I_{+}$ $=\hbar\begin{pmatrix}0&-\tfrac{\sqrt{6}}{3}&\tfrac{2\sqrt{3}}{3}&0&0&0\\ 0&0&0&\tfrac{4}{3}&\tfrac{\sqrt{2}}{3}&0\\ 0&0&0&-\tfrac{\sqrt{2}}{3}&\tfrac{4}{3}&0\\ 0&0&0&0&0&\tfrac{\sqrt{6}}{3}\\ 0&0&0&0&0&\tfrac{2\sqrt{3}}{3}\\ 0&0&0&0&0&0\end{pmatrix}\,$, (35)
$I_{3}$ $=\hbar\begin{pmatrix}1&0&0&0&0&0\\ 0&\tfrac{2}{3}&\tfrac{\sqrt{2}}{3}&0&0&0\\ 0&\tfrac{\sqrt{2}}{3}&\tfrac{1}{3}&0&0&0\\ 0&0&0&-\tfrac{2}{3}&\tfrac{\sqrt{2}}{3}&0\\ 0&0&0&\tfrac{\sqrt{2}}{3}&-\tfrac{1}{3}&0\\ 0&0&0&0&0&-1\end{pmatrix}\,$, (36)
$S_{\mathrm{h}+}$ $=\hbar\begin{pmatrix}0&\tfrac{\sqrt{6}}{3}&\tfrac{\sqrt{3}}{3}&0&0&0\\ 0&0&0&-\tfrac{1}{3}&-\tfrac{\sqrt{2}}{3}&0\\ 0&0&0&\tfrac{\sqrt{2}}{3}&\tfrac{2}{3}&0\\ 0&0&0&0&0&-\tfrac{\sqrt{6}}{3}\\ 0&0&0&0&0&\tfrac{\sqrt{3}}{3}\\ 0&0&0&0&0&0\end{pmatrix}\,$, (37)
$S_{\mathrm{h}3}$ $=\hbar\begin{pmatrix}\tfrac{1}{2}&0&0&0&0&0\\ 0&-\tfrac{1}{6}&-\tfrac{\sqrt{2}}{3}&0&0&0\\ 0&-\tfrac{\sqrt{2}}{3}&\tfrac{1}{6}&0&0&0\\ 0&0&0&\tfrac{1}{6}&-\tfrac{\sqrt{2}}{3}&0\\ 0&0&0&-\tfrac{\sqrt{2}}{3}&-\tfrac{1}{6}&0\\ 0&0&0&0&0&-\tfrac{1}{2}\end{pmatrix}\,$. (38)
The matrix elements of the annihilation operators are given by the conjugate transpose of the creation operator matrices, i.e., $I_{-}=I_{+}^{\dagger}$ and $S_{\mathrm{h}-}=S_{\mathrm{h}+}^{\dagger}$. The matrix elements of products of spin operators can be derived by multiplication of the corresponding spin matrices. Using $\bm{I}\cdot\bm{S}_{\mathrm{h}}=\frac{1}{2}(\bm{J}^{2}-\bm{I}^{2}-\bm{S}_{\mathrm{h}}^{2})$, its matrix elements are
$\langle J^{\prime}m_{J}^{\prime}|\bm{I}\cdot\bm{S}_{\mathrm{h}}|Jm_{J}\rangle=\frac{1}{2}\left(J(J+1)-\frac{11}{4}\right)\delta_{J^{\prime}J}\delta_{m_{J}^{\prime}m_{J}}\,.$ (39)